According to a Pew Research Center study, in May 2011, 30% of all American adults had a smart phone (one which the user can use to read email and surf the Internet). A communications professor at a university believes this percentage is higher among community college students. She selects 422 community college students at random and finds that 140 of them have a smart phone. Then in testing the hypotheses:

H 0 : p = 0.3

versus

H a : p > 0.3,

what is the test statistic?
Answer #1

P = X / n = 140/422 = 0.3318


Test Statistic :-
Z = ( P - P0 ) / ( √((P0 * q0)/n)
Z = ( 0.3318 - 0.3 ) / ( √(( 0.3 * 0.7) /422))
Z = 1.4235 ≈ 1.42